(20%) Problem 3: Two small balls with charges q,-37 μC and q =-12.2 μC are separated by a distance d= 5.0 cm. Calculate the clectric flux passing through the following three surfaces 33% Part (a) A spherical surface of radius 2.7 cm centered on the first charge. Grade Deduc Potent cosO cotanO asin0 acos(0 Subm Attem (4% p detail sin tan atan0 acotan) sinh0 cosh)tan cotanho Degrees Radians Submit I give up! Hints: 0 for a 0% deduction. Hints remaining: 0 Feedback: 1% deduction per feedback 33% Part (b) A spherical surface of radius 14 cm centered on the second charge 33% Part (c) A spherical surface of radius 1.1 cm centered at the midpoint of a line connecting them
Answer #1

using Gauss law

Flux = charge enclosed / epsilon_not

a)

only q1 encloses, so

F = 3.7*10^-6/(8.85*10^-12)

F = 4.181*10^5 Nm^2 /C

======

b)

since both charges enclosed by thegiven surface

F = ( 3.7 - 12.2)*10^-6 / (8.85*10^-12)

F = - 9.6*10^5 Nm^2/C

=======

c)

Since no charge enclosed

F = 0 Nm^2/C

=======

Comment in case any doubt, will reply for sure.. Goodluck